OpenStudy (anonymous):

Triple Variable System Help? 4x-y+z=0 -2x+2y+3z=3 -x+3y-2z=-19

OpenStudy (baldymcgee6):

I would use elimination by multiplying eqn two by 2, and eqn 2 by 4, then make your x's cancel

OpenStudy (anonymous):

so -4x-4y+3z=-3 after multiplying equation 2 by 2?

OpenStudy (anonymous):

i mean -4x-4y+6z=-6

OpenStudy (baldymcgee6):

2*(-2x+2y+3z=3) -4x+4y+6z = 6

OpenStudy (anonymous):

Wow I wrote the problem down wrong in so many ways :p oh well. Alright, so now -4x+4y+6z = 6 into -x+3y-2z=-19 and get rid of the x, yeah?

OpenStudy (baldymcgee6):

No, you have to multiply the 3rd eqn by 4

OpenStudy (anonymous):

I came up with 4x+12y-12z=-114